# Modelling conductive thermal transport in three-dimensional fibrous media with fiber-to-fiber contacts

### v1

@@ IV Conclusion
In summary, we have numerically generated 3D networks representative of fibrous insulation materials such as glasswool, and we have developed an original and efficient nodal simulation approach to analyze their steady-state solid thermal conductivity, integrating the effect of fiber-to-fiber thermal contact resistances. We have investigated the effects of fiber length, diameter, volume fraction and angle, and we have demonstrated that the evolution of solid conductivity in 3D networks is governed by a master curve that depends on a characteristic ratio of thermal resistances, which is an innovative way of describing the behavior of this quantity. This approach also constitutes an exhaustive validation in the 3D case of the underlying theoretical model, which had previously been numerically validated in the specific case of 2D carbon nanotube networks [48, 46, 47]; a first comparison with 3D simulation results having been presented only very recently [49], but for more limited ranges of fiber geometric parameters. In particular, a distinction of this present work is the consideration of various fiber orientations, i.e., the investigation of anisotropic systems, which is particularly important when considering insulation materials like glasswool. In this material, fibers also exhibit relatively low aspect ratios and low densities: we have shown that in these cases, a modified theoretical model is required to take into account the influence of correlations between fiber temperatures. We have in particular introduced a simple form of this new model, applicable to a wide range of fibrous materials, highlighting the significance of the average number of contacts per fiber as a key geometric parameter to describe the degree of connectivity of the networks. This model, which enables predictions of solid conductivity based solely on morphological parameters, represents a significant advancement in enhancing the accuracy of thermal performance predictions for insulation materials.
@@
All the conclusions drawn in this study also apply to the prediction of the electrical conductivity of 3D fibrous materials due to the thermal-electric analogy [66]. Therefore, this work paves the way for further investigating the conduction of heat or electricity through the solid structure in fibrous materials with more complex properties. We plan to extend the use of the nodal method to assemblies of non-straight fibers, which could be generated using an approach similar to that of Kallel et al. [27]. A lower solid conductivity is expected when introducing fiber curviness, as the effective length between the ends of fibers will be reduced, thus decreasing the number of conduction paths [40, 67]. For a more accurate representation of real fibrous samples, the variability of fiber parameters in real samples should also be accounted for, using distributions in the model, as well as the correlations existing between them. The presumed link between contact resistance and fiber diameter was indeed already discussed as an interesting next inclusion in the model, but the link between fiber length and diameter could also be explored. Finally, we plan to integrate in the model presented here the results of future experimental measurements of contact resistance between two glass fibers, in order to provide insight about the extent of the impact of contacts in real glasswool samples.
@@ Acknowledgements.
The authors are grateful to Jean-Marie Chassot (Institut Langevin) who recorded the optical coherence tomography image of a glasswool mat shown in the supplementary material, using LLTech commercial system. This work pertains to the French government program ”France2030” (EUR INTREE and LABEX INTERACTIFS). This work is supported by the ANRT (Association nationale de la recherche et de la technologie) thanks to a CIFRE fellowship granted to C. Gaunand. This work was also supported by the ”Investissements d’Avenir” program launched by the French Government (Labex WiFi).
@@ Appendix A Resolution of the linear thermal circuit system using a nodal method
The incidence matrix of a graph with $n$ nodes and $m$ branches $\boldsymbol{A}=(a_{fg})_{1\leq f\leq m,1\leq g\leq n}$ is defined by [68]:
$a_{fg}=\left\{\begin{array}[]{ll}0&\text{ \footnotesize if the branch $\it{f}$is not connected to the node $\it{g}$}\\ +1&\text{\footnotesize if the flow through the branch $\it{f}$ enters the node$\it{g}$}\\ -1&\text{\footnotesize if the flow through the branch $\it{f}$ leaves the node$\it{g}$.}\end{array}\right.$ (9)
The rows of the matrix $\boldsymbol{A}$ correspond to the branches, and the columns correspond to the nodes. As the heat fluxes through branches have defined directions, the network is a directed graph, hence an oriented incidence matrix.
@@
where $R_{f}$ is the thermal resistance associated with the branch $f$. If the branch represents a fiber, its resistance is an internal resistance: $R_{f}=\frac{4l_{f}}{k_{fib}\pi d^{2}}$ where $l_{f}$ is the length of the branch (i.e., the distance between the two associated nodes). If it represents a contact, its resistance is a contact resistance: $R_{f}=R_{k}$ where $R_{k}$ is a value we impose for all contacts in the assembly. The introduction of the contact resistance as an effective value is chosen due to the soft-core approach used in this numerical model: there is no dependence on the individual properties of the contacts.
@@
For any branch $f$, the temperature drop is $e_{f}=t_{g-1}-t_{g}+b_{f}$, where $t_{g}$ and $t_{g-1}$ are the temperatures of the connected nodes and $b_{f}$ the possible temperature source term. The heat flux through the branch is linked to this temperature difference and to the thermal resistance associated to the branch by the Fourier’s Law: $q_{f}=e_{f}/R_{f}$. Expressed for all branches in matrix form, these two equations become:
@@
where $\boldsymbol{e}$ is the $m$-dimensional vector of temperature drops over the thermal resistances. Combining Eqs. (11) and (12) gives:
@@
For any node $g$, the balance of heat rates can be expressed as $\sum_{g}q_{g}+s_{g}=0$, where $\sum_{g}q_{g}$ is the algebraic sum of heat fluxes entering the node and $s_{g}$ is a possible heat flow source. Expressed for all nodes in matrix form, this equation becomes:
@@
Combining the two constitutive equations for conductive heat transfer (A5) and (A6), we obtain:
@@
In the network configuration described above, there is no heat flux source (i.e., $\boldsymbol{s}=\boldsymbol{0}$), only temperature sources at the top and bottom nodes: they are associated to the corresponding branches, to include them in the vector $\boldsymbol{b}$. Eq. (15) is solved in order to get the vector of temperatures and that of heat fluxes.
@@ Appendix B Derivation of the low-connectivity model for solid conduction in fibrous networks
This model is built on a similar basis than that introduced by Volkov and Zhigilei [46] and Zhao et al. [48] for 2D and 3D dense networks, and that presented later by Volkov and Zhigilei [47] for 2D semi-dilute networks. In this proposed version, a corrective prefactor is introduced to accurately represent 3D fibrous networks with low connectivity, i.e., low density or low aspect ratio. In this case, the temperature gradients in the fibers can vary significantly, largely influenced by their proximity to other fibers.
In the guarded hot plate configuration, Fourier’s law gives the steady state heat flux in the vertical direction $Q$ as a function of the thermal conductivity of the material, the established temperature gradient in the vertical direction $\nabla T_{z}$, and the horizontal cross-sectional area $A_{z}=L^{2}$. As the thermal conductivity only accounts for conduction through the solid phase, Fourier’s law directly gives the solid conductivity: $Q=-k_{cond,solid}A_{z}\nabla T_{z}$. This also implies that $Q$ is the algebraic sum of all heat flows through the fibers that cross the section $A_{z}$, which means that it can be expressed by an averaging approach: $Q=-A_{z}\langle n_{z}\rangle\langle Q_{i}\rangle$ where $\langle Q_{i}\rangle$ is the average flux crossing one fiber (with $Q_{i}\geq 0$ if heat flows in the fiber from the end closer to the hot plate to the other end; this fiber direction will be referred to as $w_{i}$), and $\langle n_{z}\rangle$ is the average areal number density of fibers through a horizontal section. Note that $\langle Q_{i}\rangle\geq 0$ on average, which makes $Q$ a negative quantity, consistently with the coordinate system chosen in Fig. 1(a) and as represented in Fig. 3. The solid conductivity can therefore be given by:
@@
At the fiber scale, the flux $Q_{i}$ can be expressed by a local Fourier’s law: $Q_{i}=-k_{fib}A_{fib}\frac{dT_{i}}{dw_{i}}$ with $\frac{dT_{i}}{dw_{i}}$ the temperature gradient in the direction $w_{i}$ of the fiber $i$, $k_{fib}$ the intrinsic fiber conductivity and $A_{fib}=\pi d^{2}/4$ the fiber cross-section. The conductivity and the cross-section being held constant for all fibers in the network, an average form can be written as follows:
@@
This approach relies on the assumption of a constant average temperature gradient, that takes into account the contribution of fibers conveying heat flux in the opposite direction to the global steady-state heat flux established in the sample. Introducing the average cosine of the polar angle $\theta$, this temperature gradient can be expressed along the $z$-direction by:
@@
Hence, the temperature of fiber $i$ at a given $z$-coordinate is given by the following decomposition:
@@
where $T_{b}$ is the source temperature at $z=0$, $z_{c_{i}}$ is the $z$-coordinate of the center of the fiber, and $\delta T_{i}$ is a corrective term accounting for the deviation of the real temperature distribution along fiber $i$ from the linear one. In particular, this term can represent the contribution of correlations existing between the temperatures of two fibers in contact. For two fibers $i$ and $j$ in contact, at $z=z_{ij}$ (the position of the contact on fiber $i$) and $z=z_{ji}$ (the position of the contact on fiber $j$), respectively, Eq. (19) gives:
@@
As the values chosen in this study for fiber diameter, that acts as the distance threshold for contact detection, are about two orders of magnitude smaller than fiber length, it is reasonable to assume that $z_{ij}\approx z_{ji}$. Then, the temperature difference at a junction is given by:
@@
where $\delta T_{ij}=\delta T_{j}(z_{ji})-\delta T_{i}(z_{ij})$ represents the correction on the estimated temperature drop due to correlations. Considering heat transfer will occur from the fiber whose center is closer to the hot plate to the other one, and introducing $\langle H\rangle=\lvert z_{c_{j}}-z_{c_{i}}\rvert$ the average $z$-distance between two fibers in contact, averaging Eq. (20) results in:
@@
On average, one can assume that the flux through one fiber $\langle Q_{i}\rangle$ flows in or out through half of the contacts, as each contact corresponds to two fibers: $\langle Q_{i}\rangle=\langle\frac{1}{2}\sum_{j}\lvert Q_{ij}\rvert\rangle$, where $\lvert Q_{ij}\rvert$ is the net interface flux between fibers $i$ and $j$, governed by the value of the contact resistance: $\lvert Q_{ij}\rvert=\frac{\lvert T_{j}-T_{i}\rvert}{R_{k}}$. This flow balance can be therefore written as:
@@
where $\langle N_{c}\rangle$ is the average number of contacts per fiber. Using Eqs. (17), (18) and (22), the following relationship between the average temperature gradient along fibers in the $z$-direction and the average temperature drop at a junction is obtained:
@@
Combining Eq. (16) and Eq. (22), we have: $k_{cond,solid}=\frac{\langle n_{z}\rangle\langle N_{c}\rangle\langle\lvert\Delta T_{ij}\rvert\rangle}{2R_{k}\nabla T_{z}}$. This expression for the solid conductivity can be inserted in Eq. (24), which gives its evolution as a function of $\nabla T_{z}$, $\langle\delta T_{ij}\rangle$ and the geometric parameters of the network:
@@
Using the notations introduced in section III.2, $r=\frac{2l\langle\lvert\cos\theta\rvert\rangle}{\langle H\rangle\langle N_{c}\rangle}\frac{R_{k}}{R_{fib}}=\frac{2R_{k}\langle\lvert\cos\theta\rvert\rangle k_{fib}A_{fib}}{\langle H\rangle\langle N_{c}\rangle}$ and $k_{0,th}=k_{fib}A_{fib}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$, Eq. (25) can be written as:

### v2

@@ IV Conclusion
In summary, we have numerically generated 3D networks representative of fibrous insulation materials such as glasswool, and we have developed an original and efficient nodal simulation approach to analyze their steady-state solid thermal conductivity, integrating the effect of fiber-to-fiber thermal contact resistances. We have investigated the effects of fiber length, diameter, volume fraction and angle, and we have demonstrated that the evolution of solid conductivity in 3D networks is governed by a master curve that depends on a characteristic ratio of thermal resistances, which is an innovative way of describing the behavior of this quantity. This approach also constitutes an exhaustive validation in the 3D case of the underlying theoretical model, which had previously been numerically validated in the specific case of 2D carbon nanotube networks [48, 46, 47]; a first comparison with 3D simulation results was presented only recently [49], but for more limited ranges of fiber geometric parameters. In particular, a distinction of this present work is the consideration of various fiber orientations, i.e., the investigation of anisotropic systems, which is particularly important when considering insulation materials like glasswool. In this material, fibers also exhibit relatively low aspect ratios and low densities: we have shown that in these cases a modified theoretical model is required to take into account the influence of correlations between fiber temperatures. We have in particular introduced a simple form of this model, applicable to a wide range of fibrous materials, highlighting the significance of the average number of contacts per fiber as a key geometric parameter to describe the degree of connectivity of the networks. This model, which enables predictions of solid conductivity based solely on morphological parameters, represents a significant advancement in enhancing the accuracy of thermal performance predictions for insulation materials.
@@
All the conclusions drawn in this study also apply to the prediction of the electrical conductivity of 3D fibrous materials due to the thermal-electric analogy [66]. Therefore, this work paves the way for further investigation of the conduction of heat or electricity through the solid structure in fibrous materials with more complex properties. We plan to extend the use of the nodal method to assemblies of nonstraight fibers, which could be generated using an approach similar to that of Kallel et al. [27]. A lower solid conductivity is expected when introducing fiber curviness, as the effective length between the ends of fibers will be reduced, thus decreasing the number of conduction paths [40, 67]. For a more accurate representation of real fibrous samples, the variability of fiber parameters in real samples should also be accounted for, using distributions in the model, as well as the correlations existing between them. The presumed link between contact resistance and fiber diameter was indeed already discussed as an interesting next inclusion in the model, but the link between fiber length and diameter could also be explored. Finally, we plan to integrate in the model presented here the results of future experimental measurements of contact resistance between two glass fibers, in order to provide insight into the extent of the impact of contacts in real glasswool samples.
@@ Acknowledgements.
The authors are grateful to Jean-Marie Chassot (Institut Langevin), who recorded the optical coherence tomography image of a glasswool mat shown in the supplementary material, using LLTech commercial system. This work pertains to the French government program ”France2030” (EUR INTREE and LABEX INTERACTIFS). This work is supported by the ANRT (Association nationale de la recherche et de la technologie) thanks to a CIFRE fellowship granted to C.G. This work is also supported by the ”Investissements d’Avenir” program launched by the French Government (Labex WiFi).
@@ Appendix A Resolution of the linear thermal circuit system using a nodal method
The incidence matrix of a graph with $n$ nodes and $m$ branches, $\boldsymbol{A}=(a_{fg})_{1\leq f\leq m,1\leq g\leq n}$ is defined by [68]
$a_{fg}=\left\{\begin{array}[]{ll}0&\text{ \footnotesize if branch $\it{f}$ is not connected to node $\it{g}$,}\\ +1&\text{\footnotesize if the flow through branch $\it{f}$ enters node $\it{g}$,}\\ -1&\text{\footnotesize if the flow through branch $\it{f}$ leaves node $\it{g}$.}\end{array}\right.$ (9)
The rows of matrix $\boldsymbol{A}$ correspond to the branches, and the columns correspond to the nodes. As the heat fluxes through branches have defined directions, the network is a directed graph, and hence an oriented incidence matrix.
@@
with $R_{f}$ is the thermal resistance associated with branch $f$. If the branch represents a fiber, its resistance is an internal resistance: $R_{f}=\frac{4l_{f}}{k_{fib}\pi d^{2}}$ with $l_{f}$ the length of the branch (i.e., the distance between the two associated nodes). If it represents a contact, its resistance is a contact resistance: $R_{f}=R_{k}$ with $R_{k}$ a value we impose for all contacts in the assembly. The introduction of the contact resistance as an effective value is chosen due to the soft-core approach used in this numerical model: there is no dependence on the individual properties of the contacts.
@@
For any branch $f$, the temperature drop is $e_{f}=t_{g-1}-t_{g}+b_{f}$, where $t_{g}$ and $t_{g-1}$ are the temperatures of the connected nodes and $b_{f}$ is the possible temperature source term. The heat flux through the branch is linked to this temperature difference and to the thermal resistance associated with the branch by Fourier’s Law: $q_{f}=e_{f}/R_{f}$. Expressed for all branches in matrix form, these two equations become
@@
where $\boldsymbol{e}$ is the $m$-dimensional vector of temperature drops over the thermal resistances. Combining Eqs. (11) and (12) gives
@@
For any node $g$, the balance of heat rates can be expressed as $\sum_{g}q_{g}+s_{g}=0$, where $\sum_{g}q_{g}$ is the algebraic sum of heat fluxes entering the node and $s_{g}$ is a possible heat flow source. Expressed for all nodes in matrix form, this equation becomes
@@
Combining the two constitutive equations for conductive heat transfer [Eqs. (A5) and (A6)], we obtain
@@
In the network configuration described above, there is no heat flux source (i.e., $\boldsymbol{s}=\boldsymbol{0}$), only temperature sources at the top and bottom nodes: they are associated with the corresponding branches, to include them in vector $\boldsymbol{b}$. Equation (15) is solved in order to get the vector of temperatures and that of heat fluxes.
@@ Appendix B Derivation of the low-connectivity model for solid conduction in fibrous networks
This model is built on a similar basis to that introduced by Volkov and Zhigilei [46] and Zhao et al. [48] for 2D and 3D dense networks, and that presented later by Volkov and Zhigilei [47] for 2D semidilute networks. In this proposed version, a corrective prefactor is introduced to accurately represent 3D fibrous networks with low connectivity, i.e., low density or low aspect ratio. In this case, the temperature gradients in the fibers can vary significantly, largely influenced by their proximity to other fibers.
In the guarded hot plate configuration, Fourier’s law gives the steady-state heat flux in the vertical direction $Q$ as a function of the thermal conductivity of the material, the established temperature gradient in the vertical direction $\nabla T_{z}$, and the horizontal cross-sectional area $A_{z}=L^{2}$. As the thermal conductivity only accounts for conduction through the solid phase, Fourier’s law directly gives the solid conductivity: $Q=-k_{cond,solid}A_{z}\nabla T_{z}$. This also implies that $Q$ is the algebraic sum of all heat flows through the fibers that cross the section $A_{z}$, which means that it can be expressed by an averaging approach, i.e., $Q=-A_{z}\langle n_{z}\rangle\langle Q_{i}\rangle$, where $\langle Q_{i}\rangle$ is the average flux crossing one fiber (with $Q_{i}\geq 0$ if heat flows in the fiber from the end closer to the hot plate to the other end; this fiber direction will be referred to as $w_{i}$), and $\langle n_{z}\rangle$ is the average areal number density of fibers through a horizontal section. Note that $\langle Q_{i}\rangle\geq 0$ on average, which makes $Q$ a negative quantity, consistent with the coordinate system chosen in Fig. 1(a) and as represented in Fig. 3. The solid conductivity can therefore be given by
@@
At the fiber scale, flux $Q_{i}$ can be expressed by a local Fourier’s law: $Q_{i}=-k_{fib}A_{fib}\frac{dT_{i}}{dw_{i}}$ with $\frac{dT_{i}}{dw_{i}}$ the temperature gradient in direction $w_{i}$ of fiber $i$, $k_{fib}$ the intrinsic fiber conductivity, and $A_{fib}=\pi d^{2}/4$ the fiber cross section. With the conductivity and cross section held constant for all fibers in the network, an average form can be written as
@@
This approach relies on the assumption of a constant average temperature gradient, which takes into account the contribution of fibers conveying heat flux in the opposite direction to the global steady-state heat flux established in the sample. Introducing the average cosine of the polar angle $\theta$, this temperature gradient can be expressed along the $z$-direction as
@@
Hence, the temperature of fiber $i$ at a given $z$ coordinate is given by the decomposition
@@
where $T_{b}$ is the source temperature at $z=0$, $z_{c_{i}}$ is the $z$ coordinate of the center of the fiber, and $\delta T_{i}$ is a corrective term accounting for the deviation of the real temperature distribution along fiber $i$ from the linear one. In particular, this term can represent the contribution of correlations existing between the temperatures of two fibers in contact. For two fibers $i$ and $j$ in contact, at $z=z_{ij}$ (the position of the contact on fiber $i$) and $z=z_{ji}$ (the position of the contact on fiber $j$), respectively, Eq. (19) gives
@@
As the values chosen in this study for fiber diameter, which acts as the distance threshold for contact detection, are about 2 orders of magnitude smaller than the fiber length, it is reasonable to assume that $z_{ij}\approx z_{ji}$. Then, the temperature difference at a junction is given by
@@
where $\delta T_{ij}=\delta T_{j}(z_{ji})-\delta T_{i}(z_{ij})$ represents the correction on the estimated temperature drop due to correlations. Considering heat transfer will occur from the fiber whose center is closer to the hot plate to the other one, and introducing $\langle H\rangle=\lvert z_{c_{j}}-z_{c_{i}}\rvert$ as the average $z$ distance between two fibers in contact, averaging Eq. (20) results in
@@
On average, one can assume that the flux through one fiber $\langle Q_{i}\rangle$ flows in or out through half of the contacts, as each contact corresponds to two fibers: $\langle Q_{i}\rangle=\langle\frac{1}{2}\sum_{j}\lvert Q_{ij}\rvert\rangle$, with $\lvert Q_{ij}\rvert$ the net interface flux between fibers $i$ and $j$, governed by the value of the contact resistance, i.e., $\lvert Q_{ij}\rvert=\frac{\lvert T_{j}-T_{i}\rvert}{R_{k}}$. This flow balance can be therefore written as
@@
where $\langle N_{c}\rangle$ is the average number of contacts per fiber. Using Eqs. (17), (18) and (22), the following relationship between the average temperature gradient along fibers in the $z$ direction and the average temperature drop at a junction is obtained:
@@
Combining Eq. (16) and Eq. (22), we have $k_{cond,solid}=\frac{\langle n_{z}\rangle\langle N_{c}\rangle\langle\lvert\Delta T_{ij}\rvert\rangle}{2R_{k}\nabla T_{z}}$. This expression for the solid conductivity can be inserted into Eq. (24), which gives its evolution as a function of $\nabla T_{z}$, $\langle\delta T_{ij}\rangle$, and the geometric parameters of the network:
@@
Using the notation introduced in Sec. III.2, $r=\frac{2l\langle\lvert\cos\theta\rvert\rangle}{\langle H\rangle\langle N_{c}\rangle}\frac{R_{k}}{R_{fib}}=\frac{2R_{k}\langle\lvert\cos\theta\rvert\rangle k_{fib}A_{fib}}{\langle H\rangle\langle N_{c}\rangle}$ and $k_{0,th}=k_{fib}A_{fib}\langle n_{z}\rangle\langle\lvert\cos\theta\rvert\rangle$, Eq. (25) can be written as